OpenStudy (emogirlatemoooocow):

Bruno will spin the arrow on the spinner 3 times. What is the probability that the arrow will stop on A, then B, then C?

OpenStudy (emogirlatemoooocow):

the wheel

OpenStudy (emogirlatemoooocow):

A.\[\frac{ 3 }{ 5}\]

OpenStudy (emogirlatemoooocow):

B. \[\frac{ 1 }{ 15 }\]

OpenStudy (emogirlatemoooocow):

C.\[\frac{ 1 }{ 125 }\]

OpenStudy (emogirlatemoooocow):

D.\[\frac{ 3 }{ 125}\]

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

c

OpenStudy (emogirlatemoooocow):

thanks :)

OpenStudy (anonymous):

np